OpenStudy (konradzuse):

Integration by parts. integral of 7x/(e^x) dx

OpenStudy (konradzuse):

Changing it to 7x * e^(-x) dx

OpenStudy (konradzuse):

u = 7x 1/7 du = dx dv = e^(-x) v = -e^(-x)

OpenStudy (konradzuse):

\[7x *-e^{-x} +\frac{1}{7} \int\limits e^{-x} dx\]

OpenStudy (konradzuse):

\[7x *-e^{-x} -\frac{1}{7} e^{-x} +c?\]

OpenStudy (konradzuse):

Apparently it's not right....hmm damnit....

OpenStudy (ash2326):

let me try

OpenStudy (konradzuse):

It says that it may not need to be used by integration by parts, but I think this has to be.

OpenStudy (ash2326):

We have \[\int 7x\times e^{-x} dx\] We do need to use by parts, can't be done without it.

OpenStudy (konradzuse):

yup.

OpenStudy (konradzuse):

Check what I did above.... I'm not sure whereI messed up...

OpenStudy (ash2326):

\[\int 7x\times e^{-x} dx\] We know by parts \[\int u v=u\int v-\int (du \int v)\] \[\int 7x \times e^{-x} dx= 7x\int e^{-x} dx-\int (\frac{d}{dx} 7x \int e^{-x} dx)dx\] we get \[-7xe^{-x}-(\int 7\times -e^{-x} dx)\] \[-7xe^{-x} +\int 7e^{-x} dx\] we get \[-7xe^{-x}-7e^{-x} +c\]

OpenStudy (ash2326):

Do you get this?

OpenStudy (konradzuse):

Oh I fluttered up by doing the 1/7 huh I think I had this konfused before as well.

OpenStudy (konradzuse):

because we aren't switching to the u "world" we are staying with x....

OpenStudy (ash2326):

oops, do you get it now?

OpenStudy (konradzuse):

Yeah I keep konfusing myself with this I guess :P.

OpenStudy (konradzuse):

We keep it as du = 7dx :P.

OpenStudy (konradzuse):

not the other way around, thanks :D

OpenStudy (ash2326):

try to keep things simple, if we had \( e^{7x}\), then we'd need substitution. But still it's not necessary:)

OpenStudy (konradzuse):

:)

OpenStudy (konradzuse):

well I was just konfused because normally we are working with dx and then going to udu or w/e, but now we take du and switch back. I guess it depends on the situaiton tho....

OpenStudy (ash2326):

yeah, here substitution is not of any help.